# An algorithm for accurate and simple-looking metaphorical maps

### v1

@@ 2.1 Quality measures for Metaphorical Maps
Brinkhoff, Kriegel, Schneider, and Braun [7] defined the complexity of a polygon $P$ as a function of three quantities (see also Appendix B), that can be intuitively understood based on an example of a star-shape: (a) the frequency of $P$’s vibration, denoted by $\text{freq}(P)$, which specifies how many tips a star-shape has, (b) the amplitude of $P$’s vibration, denoted by $\text{ampl}(P)$, which specifies how long the tips of a star-shape are and (c) $P$’s convexity, denoted by $\text{conv}(P)$, the fraction of the area that is not covered within the smallest enclosing circle. The polygon complexity is then defined as $\text{compl}(P)=0.8\cdot\text{ampl}(P)\cdot\text{freq}(P)+0.2\cdot\text{conv}(P)$, which maps $P$ to values in $[0,1]$, with low values meaning low complexity.
@@ 2.2 The MS-algorithm
The MS-algorithm [15] is a typical force-directed algorithm that employs several antagonistic forces applied to the vertices of the metaphorical map, that, hopefully, at equilibrium produce a layout with good characteristics, i.e., low normalized cartographic error and low polygon complexity. The algorithm employs three forces (vertex-vertex repulsion, vertex-edge repulsion, and angular-resolution) targeted towards producing metaphorical maps of low polygon complexity and one force (air-pressure) which works towards reducing the cartographic error of the metaphorical map. Here, we only describe the air-pressure force, since our proposed algorithm modifies how this force is applied. The exact definition of the other forces can be found in Appendix C and in [15].
The normalized air pressure in region $g$, $P(g)$, exerts a force on each bounding edge $e$ based on the pressure’s magnitude and the edge’s length $\ell(e)$ in relation to the length of the entire polygonal region boundary $\text{circ}(g)$; see Appendix˜C for a formal definition of $P(g)$. Thus, the air-pressure force on edge $e$ of region $g$ was defined as $\mathbf{F}(e,g)=3P(g)\frac{\ell(e)}{\text{circ}(g)}\widehat{\mathbf{r}}$, where $\widehat{\mathbf{r}}$ is a unit vector perpendicular to $e$ directed towards outside of $g$. Force $\mathbf{F}(e,g)$ was applied to both endpoints of edge $e$.
@@ 3 Our Metaphorical Map Generation Algorithm
We performed a few experiments (refer to Appendix D) with 20-node graphs in order to determine the parameters of the algorithm, that resulted in the following values: $\textit{step}=0.02$, $\textit{iter}=1.000$ and $s_{\textit{high}}=8$. To also account for the size of the input graph, assuming that larger graphs take a longer time to converge to a good metaphorical map, we decided to set the number of iterations to $\textit{iter}=800+10n$, the value consistent with $\textit{iter}=1.000$ for $n=20$.
@@
When we applied the stiffness coefficients to the regions, we observed that they tended to create long, narrow passages, thereby reducing the visual complexity of the metaphorical map. To mitigate this effect, we introduce a new corrective weight coefficient for the pressure forces exerted on each region’s vertices, while keeping the total pressure of the region constant. These weight coefficients redistribute pressure so that vertices in narrow passages receive a larger share of the force, whereas vertices in wider parts of the region receive less. Before defining the weight coefficient, we introduce some auxiliary notation:
@@ 3.1 Dealing with non-Triangulated Graphs
We proceed to establish the barycenter visibility property as follows. We internally triangulate the graph by creating a new auxiliary vertex for each non-triangulated inner face and connecting it to every vertex of the face. Let $G^{\prime}=(V\cup V_{aux},E\cup E_{aux})$ be the resulting internally triangulated graph where $V_{aux}$ and $E_{aux}$ are the sets of auxiliary vertices and the extra edges used in the triangulation, respectively. We apply on $G^{\prime}$ Tutte’s barycentric embedding algorithm [22] that fixes the outerface on a circle and places every inner vertex on the barycenter of its neighbors, resulting in a planar straight line layout $\Gamma_{bar}(G^{\prime})$; refer to Fig.˜2.(a-b) We finish by removing the auxiliary vertices and their adjacent edges from $\Gamma_{bar}(G^{\prime})$; refer to Fig.˜2.(c). It is trivial to see that the derived layout satisfies the barycenter visibility property. Indeed, the auxiliary vertex of each non-triangulated face $f$ is placed at the barycenter of its neighbors, that is, the vertices of $f$. Therefore, in the constructed planar drawing, face $f$ has its barycenter lying inside it. Moreover, the fact that the auxiliary edges inside $f$ do not cross the boundary of $f$, ensures the visibility between the barycenter of $f$ and all of $f$’s vertices.
@@ 4 Experimental Evaluation
The analysis of question 5 can be found in Appendix E.
@@
For the generation of test data, we followed the same technique as in the evaluation of the authors of [15]. The authors of [15] generate planar straight-line drawings based on the Delaunay triangulation of a random point set. A fraction $\textit{nest}\in[0,1]$ of the points, called nesting ratio, is placed within the triangles of the initial Delaunay triangulation. It was experimentally verified in [15] that the nesting ratio has an effect on the algorithm’s performance, therefore, we also included it in our analysis. For every vertex, we generate a weight using a uniform distribution with different weight ratios of maximum to minimum weight. The above described procedure is augmented to generate non-triangulated plane graphs as follows. Following the generation of an internally triangulated plane graph, we remove a number of randomly selected internal edges in order to create non-triangulated plane graphs. We denote the ratio of removed to initial internal edges by remrem. Note that an edge is only removed if the graph remains biconnected.
@@
For the comparison of the performance of our algorithm (NEW-algorithm) against MS-algorithm, we examined how three parameters–the nesting ratio nestnest (see Fig.˜4), the weight ratio $w$ (see Fig.˜5), and the total number of nodes $n$ (see Fig.˜6)–influence both algorithms. To this end, we conducted three experiments:
@@ 5 Conclusion
Appendix A Demo application
A demo application for generation of metaphorical maps has been implemented in Javascript and is available as a web-application. In order to avoid violating the anonymity of this submission, we defer revealing the relevant url link.
Appendix B Polygon Complexity
@@
Appendix C Further details of the MS-algorithm
@@
Appendix D Setting step, iter and $s_{\textit{high}}$
@@ 8.0 $\sim$500 $\sim$500 $\sim$500 $\sim$750
Appendix E Additional evaluation considerations

### v2

@@ 1 Introduction
A demo application for generating metaphorical maps has been implemented in JavaScript using yFiles [27] and is available as a web application at http://aarg.math.ntua.gr/demos/metaphoric_maps/. All experiments were executed using a Java implementation. To allow replicability, we make the implementation and the whole graph benchmark publicly available at https://github.com/ekatsanou/metaphorical-maps.
@@ 2.1 Quality measures for Metaphorical Maps
Brinkhoff, Kriegel, Schneider, and Braun [8] defined the complexity of a polygon $P$ as a function of three quantities (see also Appendix A), that can be intuitively understood based on an example of a star-shape: (a) the frequency of $P$’s vibration, denoted by $\text{freq}(P)$, which specifies how many tips a star-shape has, (b) the amplitude of $P$’s vibration, denoted by $\text{ampl}(P)$, which specifies how long the tips of a star-shape are and (c) $P$’s convexity, denoted by $\text{conv}(P)$, the fraction of the area that is not covered within the smallest enclosing circle. The polygon complexity is then defined as $\text{compl}(P)=0.8\cdot\text{ampl}(P)\cdot\text{freq}(P)+0.2\cdot\text{conv}(P)$, which maps $P$ to values in $[0,1]$, with low values meaning low complexity.
@@ 2.2 The MS-algorithm
The MS-algorithm [16] is a typical force-directed algorithm that employs several antagonistic forces applied to the vertices of the metaphorical map, that, hopefully, at equilibrium produce a layout with good characteristics, i.e., low normalized cartographic error and low polygon complexity. The algorithm employs three forces (vertex-vertex repulsion, vertex-edge repulsion, and angular-resolution) targeted towards producing metaphorical maps of low polygon complexity and one force (air-pressure) which works towards reducing the cartographic error of the metaphorical map. Here, we only describe the air-pressure force, since our proposed algorithm modifies how this force is applied. The exact definition of the other forces can be found in Appendix B and in [16].
The normalized air pressure in region $g$, $P(g)$, exerts a force on each bounding edge $e$ based on the pressure’s magnitude and the edge’s length $\ell(e)$ in relation to the length of the entire polygonal region boundary $\text{circ}(g)$; see Appendix˜B for a formal definition of $P(g)$. Thus, the air-pressure force on edge $e$ of region $g$ was defined as $\mathbf{F}(e,g)=3P(g)\frac{\ell(e)}{\text{circ}(g)}\widehat{\mathbf{r}}$, where $\widehat{\mathbf{r}}$ is a unit vector perpendicular to $e$ directed towards outside of $g$. Force $\mathbf{F}(e,g)$ was applied to both endpoints of edge $e$.
@@ 3 Our Metaphorical Map Generation Algorithm
We performed a few experiments (refer to Appendix C) with 20-node graphs in order to determine the parameters of the algorithm, that resulted in the following values: $\textit{step}=0.02$, $\textit{iter}=1.000$ and $s_{\textit{high}}=8$. To also account for the size of the input graph, assuming that larger graphs take a longer time to converge to a good metaphorical map, we decided to set the number of iterations to $\textit{iter}=800+10n$, the value consistent with $\textit{iter}=1.000$ for $n=20$.
@@
When we applied the stiffness coefficients to the regions, we observed that they tended to create long, narrow passages, thereby reducing the visual complexity of the metaphorical map. To mitigate this effect, we introduce a new corrective weight coefficient for the pressure forces exerted on each regionÔÇÖs vertices, while keeping the total pressure of the region constant. These weight coefficients redistribute pressure so that vertices in narrow passages receive a larger share of the force, whereas vertices in wider parts of the region receive less. Before defining the weight coefficient, we introduce some auxiliary notation:
@@ 3.1 Dealing with non-Triangulated Graphs
We proceed to establish the barycenter visibility property as follows. We internally triangulate the graph by creating a new auxiliary vertex for each non-triangulated inner face and connecting it to every vertex of the face. The use of auxiliary vertices for mesh generation, often called Steiner points, is a common technique to transform general grids into high-quality triangulations (see [5, 18]). Let $G^{\prime}=(V\cup V_{aux},E\cup E_{aux})$ be the resulting internally triangulated graph where $V_{aux}$ and $E_{aux}$ are the sets of auxiliary vertices and the extra edges used in the triangulation, respectively. We apply on $G^{\prime}$ Tutte’s barycentric embedding algorithm [25] that fixes the outerface on a circle and places every inner vertex on the barycenter of its neighbors, resulting in a planar straight line layout $\Gamma_{bar}(G^{\prime})$; refer to Fig.˜2.(a-b) We finish by removing the auxiliary vertices and their adjacent edges from $\Gamma_{bar}(G^{\prime})$; refer to Fig.˜2.(c). It is trivial to see that the derived layout satisfies the barycenter visibility property. Indeed, the auxiliary vertex of each non-triangulated face $f$ is placed at the barycenter of its neighbors, that is, the vertices of $f$. Therefore, in the constructed planar drawing, face $f$ has its barycenter lying inside it. Moreover, the fact that the auxiliary edges inside $f$ do not cross the boundary of $f$, ensures the visibility between the barycenter of $f$ and all of $f$’s vertices.
@@ 4 Experimental Evaluation
The analysis of question 5 can be found in Appendix D.
@@
For the generation of test data, we followed the approach of [16] (described in detail in Algorithm 5.1 of [20] with a uniform distribution). The authors of [16] generate planar straight-line drawings based on the Delaunay triangulation of a random point set. A fraction $\textit{nest}\in[0,1]$ of the points, called nesting ratio, is placed within the triangles of the initial Delaunay triangulation. It was experimentally verified in [16] that the nesting ratio has an effect on the algorithm’s performance, therefore, we also included it in our analysis. For every vertex, we generate a weight using a uniform distribution with different weight ratios of maximum to minimum weight. The above described procedure is augmented to generate non-triangulated plane graphs as follows. Following the generation of an internally triangulated plane graph, we remove a number of randomly selected internal edges in order to create non-triangulated plane graphs. We denote the ratio of removed to initial internal edges by remrem. Note that an edge is only removed if the graph remains biconnected.
@@
For the comparison of the performance of our algorithm (NEW-algorithm) against MS-algorithm, we examined how three parametersÔÇöthe nesting ratio nestnest (see Fig.˜4), the weight ratio $w$ (see Fig.˜5), and the total number of nodes $n$ (see Fig.˜6)–influence both algorithms. To this end, we conducted three experiments:
@@ 5 Conclusion
Appendix A Polygon Complexity
@@
Appendix B Further details of the MS-algorithm
@@
Appendix C Setting step, iter and $s_{\textit{high}}$
@@ 8.0 $\sim$500 $\sim$500 $\sim$500 $\sim$750
Appendix D Additional evaluation considerations